Two angles are complementary. One angle measures 10 degree more than three times the measure of the other angle. Find the measure of each angle.
Answer by nerdybill(7384)

Two angles are complementary. One angle measures 10 degree more than three times the measure of the other angle. Find the measure of each angle.
.
if two angles are "complementary" -- the sum adds up to be 90 degrees.
.
Let x = one angle
then
3x+10 = second angle
.
x + 3x+10 = 90
4x+10 = 90
4x = 80
x = 20 degrees (one angle)
.
second angle:
3x+10 = 3(20)+10 = 60+10 = 70 degrees